# Blessing of dimension in Bayesian inference on covariance matrices

### v1

@@ 3.3 Uncertainty Quantification
We now consider uncertainty quantification of entrywise elements of the true covariance matrix $\Psi_{0}$. For $1\leq u,v\leq p,$ let ${\Psi}_{C,uv}$ and $\Psi_{0,uv}$ denote the $uv$th element of ${\Psi}_{C}$ and $\Psi_{0}$, respectively, where $\Psi_{C}$ denotes the coverage-corrected pseudo-posterior sample obtained from the CC-FABLE discussed in Section 2.2. For $1\leq u,v\leq p$, we have $\Psi_{0,uv}=\lambda_{0u}^{\top}\lambda_{0v}+\sigma_{0u}^{2}\mathbbm{1}(u=v)$, where $\mathbbm{1}$ denotes the indicator function. In what follows, we discuss results for a fixed value of the coverage-correction factor $\rho>0$ and then provide an approach to appropriately choose $\rho$.
We now let $T_{uv}=\mu_{u}^{\top}\mu_{v}+\delta_{u}^{2}\mathbbm{1}(u=v)$ be an estimator of $\Psi_{0,uv}$. Then, we have the following result approximating the pseudo-posterior distribution of ${\Psi}_{C,uv}$ by a suitable Gaussian distribution centered at $T_{uv}$ as both $n,p_{n}$ increase with $\rho>0$ fixed.
@@
Suppose Assumptions 3.13.4 hold and let $\sqrt{n}/p_{n}=o(1)$. For $1\leq u,v\leq p$, let
$\begin{split}l_{0,uv}^{2}(\rho)&=\left\{\begin{array}[]{ll}\rho^{2}(\sigma_{0v}^{2}\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2}\|\lambda_{0v}\|_{2}^{2}),&\text{for $u\neq v$},\\ 2\sigma_{0u}^{4}+4\rho^{2}\,\sigma_{0u}^{2}\|\lambda_{0u}\|_{2}^{2},&\text{for$u=v$.}\end{array}\right.\end{split}$
Then, as $n\to\infty,$
$\sup_{x\in\mathbb{R}}\,\left|{\Pi}_{C}\left\{\dfrac{\sqrt{n}({\Psi}_{C,uv}-T_{uv})}{\l_{0,uv}(\rho)}\leq x\right\}-\Phi(x)\right|\overset{P_{0}}{\to}0\quad\text{for all $1\leq u,v\leq p.$}$
Theorem 3.7 allows us to approximate the asymptotic pseudo-posterior distribution of each element of the covariance matrix after suitable centering and scaling using a Gaussian distribution with mean 00 and variance $l_{0,uv}^{2}(\rho)$. The proof of Theorem 3.7 is in Section C of the Supplementary Material. We next state a result regarding the asymptotic law of the quantity $\sqrt{n}\,(T_{uv}-\Psi_{0,uv})/\mathcal{S}_{0,uv}$, demonstrate how this result can be used to obtain asymptotic entrywise frequentist coverage for $\Psi_{0,uv}$ using CC-FABLE, and outline different choices of $\rho$. For general random variables $X_{n}$ and $X$, we denote $X_{n}$ converging in distribution to $X$ by $X_{n}\implies X.$
@@
Suppose the Assumptions of Theorem 3.7 hold. For $1\leq u,v\leq p$, let
$\begin{split}\mathcal{S}_{0,uv}^{2}&=\left\{\begin{array}[]{ll}\sigma_{0v}^{2}\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2}\|\lambda_{0v}\|_{2}^{2}+\|\lambda_{0u}\|_{2}^{2}\|\lambda_{0v}\|_{2}^{2}+(\lambda_{0u}^{\top}\lambda_{0v})^{2},&\text{for $u\neq v$},\\ 2(\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2})^{2},&\text{for $u=v$.}\end{array}\right.\end{split}$
Then, as $n\to\infty$, one has $\sqrt{n}(T_{uv}-\Psi_{0,uv})/\mathcal{S}_{0,uv}\implies N(0,1)$ for all $1\leq u,v\leq p.$
The proof of Theorem 3.8 is available in Section D of the Supplementary Material. Under Theorem 3.7, the $100(1-\alpha)\%$ asymptotic pseudo-posterior credible interval of ${\Psi}_{C,uv}$ is given by
$\mathcal{C}_{uv}(\rho)=\left[T_{uv}\pm z_{1-(\alpha/2)}\dfrac{l_{0,uv}(\rho)}{\sqrt{n}}\right],$
@@
The form of $q_{uv}(\rho)$ suggests a general data-adaptive framework to choose $\rho$. For example, we can ensure that all the entrywise asymptotic coverages are at least $1-\alpha$ by letting $\rho=\sup_{u,v}b_{uv}$ for $1\leq u\leq v\leq p$, with $b_{uv}$ defined in Section 2.2. To see why, first let $\hat{q}_{uv}(\rho)$ be the consistent estimator of $q_{uv}(\rho)$ obtained by substituting $\sigma_{0j}^{2}$ and $\lambda_{0j}$ with $\mathcal{V}_{j}^{2}$ and $\mu_{j}$, respectively, for $j=1,\ldots,p$; these quantities are defined in Section 2. Next, observe that $\hat{q}_{uv}(\rho)$ is a non-decreasing function of $\rho>0$ for all $1\leq u,v\leq p$. It is clear that $\hat{q}_{uv}(b_{uv})=1-\alpha$ for all $1\leq u\leq v\leq p.$ Thus, $\hat{q}_{uv}\left(\sup_{u^{\prime},v^{\prime}}b_{u^{\prime}v^{\prime}}\right)\geq\hat{q}_{uv}(b_{uv})=1-\alpha.$ Alternatively, we can ensure the average of the entrywise asymptotic coverages is $1-\alpha$ by numerically solving
$\dfrac{2}{p(p+1)}\displaystyle\sum_{1\leq u\leq v\leq p}\hat{q}_{uv}(\rho)=1-\alpha$ (19)
for $\rho$. We observed in simulation experiments that simply letting
$\rho=\overline{b}=\left(\sum_{1\leq u\leq v\leq p}b_{uv}\right)/\left\{\dfrac{p(p+1)}{2}\right\}$
is a close approximation to the root of (19). In practice, we thus recommend choosing $\rho=\overline{b}$ for computational convenience, as this avoids solving the nonlinear equation. This worked well in the simulations we carried out, both from the perspective of median and average asymptotic entrywise coverages. Since Lemma E.5 in Section E of the Supplementary Material ensures that the matrix $B=(b_{uv})_{1\leq u,v\leq p}$ satisfies $\|B\|_{\infty}=O_{P_{0}}(1)$, all the aforementioned estimates of $\rho$ are bounded.
@@ 4.1 Setup
In this Section, we compare the proposed CC-FABLE approach with competitors when judging their performance from the viewpoint of estimation error, uncertainty quantification (UQ), and computational efficiency. With $\Lambda_{0}=(\Lambda_{0,uv})_{1\leq u,v\leq p}$ the true matrix of loadings and $\Sigma_{0}=\mbox{diag}(\sigma_{01}^{2},\ldots,\sigma_{0p}^{2})$ the true matrix of idiosyncratic error variances, the data generating covariance matrix is given by $\Psi_{0}=\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0}$. For both the estimation error and UQ experiments, we proceed as follows. We vary the number of samples $n\in\{500,1000\}$, the number of dimensions $p\in\{1000,5000\}$, and let the true number of factors be $k=10.$ For each possible $(n,p)\in\{500,1000\}\times\{1000,5000\}$, we generate $(\Lambda_{0},\Sigma_{0})$ as
$\Lambda_{0,jl}$ $\overset{ind}{\sim}0.5\,\widetilde{\delta}_{0}+0.5\,N(0,0.5^{2})$ for all $1\leq j\leq p,1\leq l\leq k$, and
$\sigma_{0j}^{2}$ $\overset{ind}{\sim}\mathcal{U}(0.5,5),$ for all $1\leq j\leq p,$
where $\widetilde{\delta}_{0}$ denotes a point mass at 00. For each combination $(n,p)$, we use the generated $\Psi_{0}=\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0}$ and replicate the data generating process $R=100$ times. Given an estimator $\widehat{\Psi}_{0}$ of $\Psi_{0}$, we judge the estimator using the relative spectral error, defined as
$\mathcal{L}(\Psi_{0},\widehat{\Psi}_{0})=\dfrac{\|\Psi_{0}-\widehat{\Psi}_{0}\|}{\|\Psi_{0}\|},$
@@
where $\|A\|=\sup_{\|x\|_{2}=1}\|Ax\|_{2}$. When using a Bayesian method, we consider $\widehat{\Psi}_{0}$ to be the posterior mean. For a given $(n,p)$ and a particular replicate, we obtain $\mathcal{L}(\Psi_{0},\widehat{\Psi}_{0})$ given an estimator $\widehat{\Psi}_{0}$ and then average this quantity over the replicates. For the UQ experiment, we simply consider the $95\%$ entrywise credible/pseudo-credible intervals obtained and compute the frequentist coverage over the $R=100$ replicates. To ease the computational burden, we consider the coverage of a randomly chosen $100\times 100$ submatrix of $\Psi_{0}$, corresponding to the covariance of $100$ randomly chosen variables held fixed across replicates, for each $(n,p).$ Lastly, we present a runtime analysis of the different methods considered in the simulation studies. First, we compute the theoretical computational complexity of the proposed CC-FABLE method. Next, we carry out numerical experiments to compare the computational runtimes for the different approaches.
As competitors to the proposed CC-FABLE approach, we consider the multiplicative gamma shrinkage prior formulation (Bhattacharya and Dunson,, 2011) and the approach in Ročková and George, (2016) and denote them by MGSP and ROTATE, respectively, across all the simulations. The approach in Bhattacharya and Dunson, (2011) implements posterior sampling with Markov chain Monte Carlo (MCMC), yielding posterior samples of the covariance matrix, from which we extract the posterior mean and entrywise $95\%$ credible intervals. In contrast, the approach of Ročková and George, (2016) only provides a point estimate of the covariance matrix, using expectation-maximization (EM) algorithms to approximate the posterior mean corresponding to spike and slab prior distributions on the factor loadings. As a result, we do not consider ROTATE for our UQ experiments. For CC-FABLE, we consider $1000$ Monte Carlo iterates; for estimation error experiments, we use the pseudo-posterior mean, which is analytically available. For MGSP, we consider $3000$ MCMC iterates and discard the first $1000$ as burn-in. As described earlier, CC-FABLE does not require MCMC for inference, instead providing direct Monte Carlo samples. To implement MGSP, we used the infinitefactor package (Poworoznek et al.,, 2021), while code to implement ROTATE was obtained from http://veronikarock.com/FACTOR_ANALYSIS.zip. All the simulations were carried out in the R programming language (R Core Team,, 2021).
@@ 4.2 Estimation Performance
We show the results from the simulation experiments in Table 1. For all four different combinations of $(n,p)$, the performance of CC-FABLE is either the best or very close to being the best. Both CC-FABLE and MGSP perform uniformly better than the ROTATE approach across the cases. For the most challenging case with $n=500$ and $p=5000$, CC-FABLE performs noticeably better than MGSP. We believe the well-documented issues in mixing when implementing MCMC when sampling from the posterior for MGSP to be a contributing factor. Furthermore, MGSP seems to be more affected than either CC-FABLE or ROTATE when the number of dimensions increases.
Table 1: Comparison of estimation error between CC-FABLE, MGSP, and ROTATE.
Cases CC-FABLE MGSP ROTATE
$n=500,p=1000$ 0.31 0.34 0.43
$n=1000,p=1000$ 0.22 0.21 0.28
$n=500,p=5000$ 0.32 0.40 0.43
$n=1000,p=5000$ 0.24 0.25 0.29

### v3

@@ 3.3 Uncertainty Quantification
We now consider uncertainty quantification of entrywise elements of the covariance matrix. For $1\leq u,v\leq p,$ let $\widetilde{\Psi}_{uv}$ and $\Psi_{0,uv}$ denote the $(u,v)$th element of $\widetilde{\Psi}$ and $\Psi_{0}$, respectively. For $1\leq u,v\leq p$, $\Psi_{0,uv}=\lambda_{0u}^{\top}\lambda_{0v}+\sigma_{0u}^{2}\mathbbm{1}(u=v)$, where $\mathbbm{1}$ denotes the indicator function. We now discuss theoretical results for a fixed value of the coverage-correction factor $\rho>0$ and then provide an approach to appropriately choose $\rho$.
We let $T_{uv}=\mu_{u}^{\top}\mu_{v}+\delta_{u}^{2}\mathbbm{1}(u=v)$ be an estimator of $\Psi_{0,uv}$, with $\mu_{u},\delta_{u}^{2}$ for $1\leq u\leq p$ as defined in (12). Then, for any fixed $\rho\geq 1$, the following result approximates the FABLE-posterior of $\widetilde{\Psi}_{uv}$ with a suitable Gaussian distribution centered at $T_{uv}$ as both $n,p_{n}$ increase. Let $\Phi$ denote the cumulative distribution function of the $N(0,1)$ distribution.
@@
Suppose Assumptions 3.13.4 hold and fix $\rho\geq 1$. For $1\leq u,v\leq p_{n}$, let
[…]
$\text{Then, as $n\to\infty,$}\quad\sup_{x\in\mathbb{R}}\,\left|\widetilde{\Pi}\left\{\dfrac{\sqrt{n}(\widetilde{\Psi}_{uv}-T_{uv})}{l_{0,uv}(\rho)}\leq x\right\}-\Phi(x)\right|\overset{P_{0}}{\to}0\quad\text{for all $1\leq u,v\leq p_{n}.$}$
Theorem 3.8 allows us to approximate the asymptotic FABLE-posterior distribution of each element of the covariance matrix after suitable centering and scaling using a Gaussian distribution with mean 0 and variance $l_{0,uv}^{2}(\rho)$. The proof of Theorem 3.8 is in Section C of the Supplementary Material.
We next state a result regarding the asymptotic law of the quantity $\sqrt{n}\,(T_{uv}-\Psi_{0,uv})$ and illustrate how this result can be used to show asymptotic frequentist validity of entrywise credible intervals for $\Psi_{0,uv}$. We will require Assumption 3.5 together with Assumptions 3.1-3.4 that were used to establish Theorems 3.7 and 3.8. For generic random variables $X_{n}$ and $X$, we denote $X_{n}$ converging in distribution to $X$ by $X_{n}\implies X.$
@@
Suppose Assumptions 3.13.5 hold. For $1\leq u,v\leq p_{n}$, let
[…]
Then, as $n\to\infty$, one has $\sqrt{n}(T_{uv}-\Psi_{0,uv})/\mathcal{S}_{0,uv}\implies N(0,1)$ for all $1\leq u,v\leq p_{n}.$
The proof of Theorem 3.9 is available in Section D of the Supplementary Material. Under Theorem 3.8, the $100(1-\alpha)\%$ asymptotic credible interval of $\widetilde{\Psi}_{uv}$ is given by
$\mathcal{C}_{uv}(\rho)=\left[T_{uv}-z_{1-(\alpha/2)}\,\dfrac{l_{0,uv}(\rho)}{\sqrt{n}},\,T_{uv}+z_{1-(\alpha/2)}\,\dfrac{l_{0,uv}(\rho)}{\sqrt{n}}\right],$
@@
We next obtain a consistent estimator of the asymptotic coverage $q_{uv}(\rho)$ for any $\rho>0.$ This will allow us to estimate $\rho$ as outlined in Section 2.3. First, we define the quantities $\widehat{l}_{0,uv}(\cdot)$ and $\widehat{\mathcal{S}}_{0,uv}$ as plug-in estimates of $l_{0,uv}(\cdot)$ and $\mathcal{S}_{0,uv}$, respectively, obtained by replacing $\lambda_{0u}$ and $\sigma_{0u}^{2}$ by their consistent estimators $\mu_{u}$ and $\mathcal{V}_{u}^{2}$, respectively, as defined in Section 2.3. Lemma E.4 in Section E of the Supplementary Material provides a proof of their consistency. Next, we define $\widehat{q}_{uv}(\rho)$ as
$\widehat{q}_{uv}(\rho)=2\Phi\left\{z_{1-(\alpha/2)}\dfrac{\widehat{l}_{0,uv}(\rho)}{\widehat{\mathcal{S}}_{0,uv}}\right\}-1.$ (21)
Since $\widehat{l}_{0,uv}(\rho)$ is monotonically increasing in $\rho$ for $\rho\geq 1$, so is $\widehat{q}_{uv}(\rho).$ The function $\widehat{q}_{uv}(\rho)$ may now be utilized to obtain a data-adaptive estimate of $\rho$, starting from the fact that $\widehat{q}_{uv}(b_{uv})=1-\alpha$ and then following the discussion in Section 2.3. The expression of $b_{uv}$ as in (1) is obtained by solving the equation $\widehat{l}_{0,uv}(\rho)=\widehat{S}_{0,uv}$ for $\rho$, for $1\leq u\leq v\leq p$. Since Lemma E.5 in Section E of the Supplementary Material ensures that $\underset{1\leq u\leq v\leq p_{n}}{\max}b_{uv}=O_{P_{0}}(1)$, all the estimates of $\rho$ as described in Section 2.3 are bounded.
@@ 4.1 Setup
In this Section, we compare the performances of FABLE with competitors in terms of estimation error, uncertainty quantification (UQ), and computational efficiency. The competitors are as follows:
  1. (a)
    For estimation error, we compare with: (i) the multiplicative gamma shrinkage prior approach of Bhattacharya and Dunson, (2011) denoted by MGSP, (ii) the automatic rotation to sparsity approach of Ročková and George, (2016) denoted by ROTATE, (iii) the hard thresholding approach of Bickel and Levina, 2008a denoted by HT, (iv) the SCAD penalty (Fan and Li,, 2001) applied to entries of the sample covariance and then following a compromise between hard and soft thresholding as in Rothman et al., (2009), and (v) the Ledoit and Wolf, (2004) linear shrinkage estimator denoted by LW.
  2. (b)
    For UQ, we consider the Bayesian approaches FABLE and MGSP, as the rest of the approaches only provide point estimates.
  3. (c)
    Runtime analysis includes the theoretical computational complexity of FABLE, along with numerical experiments that compare computational runtimes for the different approaches.
Suppose $\Psi_{0}=\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0}$ denotes the true covariance matrix with $\Lambda_{0}=(\Lambda_{0,jl})_{1\leq j\leq p,1\leq l\leq k}$ the true loadings matrix and $\Sigma_{0}=\mbox{diag}(\sigma_{01}^{2},\ldots,\sigma_{0p}^{2})$ the true idiosyncratic error variance matrix. We vary the number of samples $n\in\{500,1000\}$, the number of dimensions $p\in\{1000,5000\}$, and let the true number of factors be $k=10.$ For each possible $(n,p)\in\{500,1000\}\times\{1000,5000\}$, we generate $(\Lambda_{0},\Sigma_{0})$ as $\Lambda_{0,jl}\overset{ind}{\sim}\pi_{0}\,\widetilde{\delta}_{0}+(1-\pi_{0})\,N(0,0.5^{2})$ for all $1\leq j\leq p,1\leq l\leq k$, and $\sigma_{0j}^{2}\overset{ind}{\sim}\mathcal{U}(0.5,5),$ for all $1\leq j\leq p,$ with $\widetilde{\delta}_{0}$ a point mass at 0. Hence, we simulate exact sparsity in the loadings, representing a challenging misspecified case for FABLE and MGSP, which use continuous shrinkage priors.
For each combination $(n,p)$, we use the generated $\Psi_{0}=\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0}$ and replicate the data generating process $R$ times. For the estimation error experiments, we take $R=50$ and consider $\pi_{0}=0.5$ (moderate sparsity) and $\pi_{0}=0.85$ (high sparsity). We assess the difficulty of each simulation setup with the average proportion of variance explained by the signal expressed in percentage, computed as $P_{av}=R_{av}/(1+R_{av}).$ Here, $R_{av}=E\|\lambda_{0j}\|_{2}^{2}/E(\sigma_{0j}^{2})$ is the average signal-to-noise ratio (SNR), where $\lambda_{0j}$ is the $j$th row of $\Lambda_{0}$. In this case, we have $R_{av}=k\,(1-\pi_{0})\,v_{0}^{2}/\bar{S}^{2}$, where $\bar{S}^{2}=0.5\times(5+0.5)=2.75$ is the average error variance and $v_{0}^{2}=0.5^{2}$ is the slab variance. For the UQ experiments, we let $R=100$ and consider $\pi_{0}=0.5$.
Next, we discuss the performance metrics for judging the estimation error and UQ experiments. Given an estimator $\widehat{\Psi}_{0}$ of $\Psi_{0}$ obtained from a particular implementation, we assess its efficacy with the relative spectral error, defined as
$\mathcal{L}(\Psi_{0},\widehat{\Psi}_{0})=\dfrac{\|\Psi_{0}-\widehat{\Psi}_{0}\|}{\|\Psi_{0}\|}.$
@@
When using a Bayesian approach, we let $\widehat{\Psi}_{0}$ be the corresponding posterior mean. For a given $(n,p)$ and a particular replicate $r=1,\ldots,R$ yielding the estimate $\widehat{\Psi}_{0}^{(r)}$, we obtain $\mathcal{L}^{(r)}(\Psi_{0},\widehat{\Psi}_{0}^{(r)})$ and proceed to report the average, $2.5\%$ quantile, and $97.5\%$ quantile of this quantity over the $R=50$ replicates. For the UQ experiments, we simply evaluate the frequentist coverage of the corresponding $95\%$ entrywise credible intervals obtained over the $R=100$ replicates. To ease the computational burden, we consider the coverage of a randomly chosen $100\times 100$ submatrix of $\Psi_{0}$, corresponding to the covariance of $100$ randomly chosen variables. For each $(n,p)$, these variable indices are held fixed across replicates.
All the methods are implemented in the R programming language (R Core Team,, 2021). Open-source code for implementing FABLE is available at https://github.com/shounakch/FABLE. To implement MGSP, we use the infinitefactor package (Poworoznek et al.,, 2025). Code used to implement the ROTATE approach was obtained from http://veronikarock.com/FACTOR_ANALYSIS.zip. To implement the methods HT, SCAD, and LW, we used the thresholdingEst, scadEst, and linearShrinkLWEst functions, respectively, from the cvCovEst package (Boileau et al.,, 2021). For FABLE, we collect $1000$ Monte Carlo samples; for estimation error experiments, we use the FABLE-posterior mean, which is explicitly available without sampling. For MGSP, we obtain $3000$ MCMC iterates and discard the first $1000$ as burn-in.
Throughout all simulations, the rank $k$ of the signal when fitting FABLE, MGSP, and ROTATE is treated as an unknown parameter and is thus estimated from the data. For the ROTATE approach, $k$ is estimated with an Indian buffet process (IBP) prior as described in Ročková and George, (2016). For the MGSP approach, $k$ is estimated within the MCMC algorithm as described in Bhattacharya and Dunson, (2011). For FABLE, we estimate $k$ following Section 2.4.1.
The ROTATE code requires the choice of a hyperparameter lambda0 that crucially affects its performance. As recommended in Section 7 of Ročková and George, (2016), we treat lambda0 as an inverse temperature parameter, and gradually increase it from $0.001$ to $18.001$ in increments of $2$. As described in their paper, the fit for a sufficiently large lambda0 approximates the maximum a posteriori (MAP) estimator of the covariance matrix. For HT, SCAD, and LW, we obtain improved performance by standardizing the data before analysis and transforming the estimated covariance back to the original units in a post-hoc step. HT, SCAD, and LW are implemented with 5-fold cross-validation by default.
@@ 4.2 Estimation Performance
Table 1: Comparison of estimation error between multiple approaches, with $50\%$ sparsity in spike-and-slab factor loadings. The “Mean” and “Range” columns show the average and $2.5\%-97.5\%$ quantiles across replicates, respectively.
$(n,p)$ $(500,1000)$ $(1000,1000)$ $(500,5000)$ $(1000,5000)$
Method Mean Range Mean Range Mean Range Mean Range
FABLE 0.32 0.28 – 0.38 0.23 0.20 – 0.27 0.33 0.30 – 0.38 0.24 0.21 – 0.28
MGSP 0.33 0.27 – 0.42 0.22 0.18 – 0.27 0.38 0.28 – 0.50 0.24 0.19 – 0.31
ROTATE 0.46 0.41 – 0.51 0.30 0.27 – 0.35 0.48 0.44 – 0.56 0.32 0.28 – 0.35
HT 0.32 0.27 – 0.39 0.22 0.20 – 0.25 0.35 0.30 – 0.40 0.23 0.21 – 0.27
SCAD 0.34 0.29 – 0.42 0.23 0.20 – 0.28 0.37 0.31 – 0.43 0.25 0.22 – 0.30
LW 0.35 0.31 – 0.39 0.25 0.21 – 0.29 0.37 0.34 – 0.42 0.26 0.23 – 0.30
Table 2: Comparison of estimation error between multiple approaches, with $85\%$ sparsity in block-diagonal factor loadings. The “Mean” and “Range” columns show the average and $2.5\%-97.5\%$ quantiles across replicates, respectively.
$(n,p)$ $(500,1000)$ $(1000,1000)$ $(500,5000)$ $(1000,5000)$
Method Mean Range Mean Range Mean Range Mean Range
FABLE 0.24 0.21 – 0.30 0.17 0.14 – 0.22 0.24 0.21 – 0.31 0.17 0.14 – 0.22
ROTATE 0.28 0.20 – 0.38 0.12 0.08 – 0.18 0.23 0.11 – 0.35 0.09 0.07 – 0.13
HT 0.17 0.13 – 0.25 0.09 0.07 – 0.13 0.17 0.13 – 0.24 0.09 0.07 – 0.13
SCAD 0.17 0.14 – 0.24 0.13 0.10 – 0.17 0.17 0.13 – 0.24 0.13 0.09 – 0.17
LW 0.26 0.19 – 0.34 0.18 0.15 – 0.21 0.26 0.20 – 0.34 0.18 0.15 – 0.22
We show results of simulation experiments for $\pi_{0}=0.5$ in Table 1. The average proportion of variance explained by the signal is $P_{av}\approx 31\%$. FABLE performs competitively for all four choices of $(n,p)$ compared to approaches that induce exact sparsity, such as ROTATE, HT, and SCAD. The performance of MGSP declines with increasing dimension more than the competitors, possibly due to mismatch between the continuous shrinkage prior and the exact sparsity in the simulated loadings. The MGSP MCMC algorithm exhibited efficient mixing, with the average effective sample size (ESS) of the $p$ idiosyncratic error variances $\sim 95\%$ of the post-burn-in samples. We also manually tuned the lambda0 parameter when implementing ROTATE and obtained estimation errors almost identical to that of FABLE for lambda0 = 2.001. However, manual tuning requires knowledge of the true covariance matrix and thus is not feasible in practice. In Section G of the Supplementary Material, we provide (i) further results for this simulation case with $85\%$ sparsity in the loadings ($\pi_{0}=0.85,\,P_{av}\approx 12\%$) and (ii) results for additional cases with smaller $n,p$ and larger $k=50$.
We also consider another case with $85\%$ sparsity in the loadings, but with a block-diagonal structure. This example is taken from Section 4 of Ročková and George, (2016). We let $k=10$ and vary $(n,p)$ as in the above case. For a given $p$, the true loading matrix has $n_{1}(p)=\lfloor(1-0.85)\,p\rfloor=\lfloor 0.15\,p\rfloor$ non-zero elements in each column and an overlap of $n_{2}(p)=\lceil 0.37\,n_{1}(p)\rceil$ non-zero entries with the successive column, with the error variances generated as before. All non-zero entries of the true loading matrix are equal to $1$. The average proportion of the variance explained by the signal is $P_{av}\approx 35\%$ for both $p=1000$ and $p=5000$. The results are in Table 2, excluding MGSP which is not competitive with the other approaches. ROTATE, HT, and SCAD outperform FABLE for most of the choices of $(n,p)$. However, FABLE remains competitive in terms of estimation error for both choices of $p$ when $n=500$. It is not surprising that ROTATE and the other approaches enforcing exact sparsity perform particularly well in the highly sparse regime; we expect FABLE and MGSP to outperform competitors when loadings are simulated under continuous light-tailed priors. However, it is notable that FABLE badly outperforms MGSP in this highly sparse case; it may be that pre-estimating the factors conveys greater robustness to a variety of true loadings structures.